esponses/submit?pos=88dep=23498325&tags-autosave question3806333 Sketch the curve, = 1 - sin(6) 5
Answer #1


* Given 8=1- sino = foption (4 correct 0=0 - 8=t- sino =1-0=1 o I 8=1 Sint = 1-1=0 o=t &r= t- sint =1-0=1 o=32 trat sint = 1+